OpenStudy (anonymous):

Given the following functions f(x) and g(x), solve (f + g)(3) and select the correct answer below: f(x) = 6x + 3 g(x) = x − 7 4 17 25 31

OpenStudy (anonymous):

@Haseeb96 @welshfella

OpenStudy (anonymous):

@Haseeb96

OpenStudy (haseeb96):

just add both these function as we add and then add like terms as I got (f+g)(x) = 7x -4 now put x =3 (f+g)(3) = 7(3) -4 = 21 -4 = 17 is your answer